### v1

# DetailCLIP: Detail-Oriented CLIP for Fine-Grained Tasks

Abstract
In this paper, we introduce DetailCLIP: A Detail-Oriented CLIP to address the limitations of contrastive learning-based vision-language models, particularly CLIP, in handling detail-oriented and fine-grained tasks like segmentation. While CLIP and its variants excel in the global alignment of image and text representations, they often struggle to capture the fine-grained details necessary for precise segmentation. To overcome these challenges, we propose a novel framework that employs patch-level comparison of self-distillation and pixel-level reconstruction losses, enhanced with an attention-based token removal mechanism. This approach selectively retains semantically relevant tokens, enabling the model to focus on the image’s critical regions aligned with the specific functions of our model, including textual information processing, patch comparison, and image reconstruction, ensuring that the model learns high-level semantics and detailed visual features. Our experiments demonstrate that DetailCLIP surpasses existing CLIP-based and traditional self-supervised learning (SSL) models in segmentation accuracy and exhibits superior generalization across diverse datasets. DetailCLIP represents a significant advancement in vision-language modeling, offering a robust solution for tasks that demand high-level semantic understanding and detailed feature extraction. https://github.com/KishoreP1/DetailCLIP.
1 Introduction
The rapid advancements in computer vision have led to the development of robust models that can understand and interpret visual data with impressive accuracy. Among these, CLIP (Contrastive Language-Image Pretraining) [38] stands out as a pioneering approach that leverages large-scale contrastive learning between images and text to create a shared embedding space. This has proven immensely successful in tasks like classification, where the model can understand and relate visual content to textual descriptions without direct supervision [51, 44, 25].
However, while CLIP excels in generalization and high-level semantic understanding, its reliance on contrastive losses presents significant challenges when adapting the model for more granular tasks like image segmentation [44, 58, 28, 48]. Fine-grained tasks like segmentation requires a comprehensive understanding of the entire scene and the precise delineation of object boundaries at the pixel level. However, the global representations learned through contrastive learning often fail to capture this level of detail, leading to suboptimal performance when using CLIP for granular tasks [21, 6, 55, 23, 37, 43].
Furthermore, traditional self-supervised learning (SSL) approaches [56, 5, 22, 57, 34, 14], which have gained popularity for their ability to learn from unlabeled data, often prove inadequate when applied to fine-grained tasks. These methods, including contrastive learning [35, 7, 41, 17] or clustering-based techniques [32], primarily focus on learning representations that are useful for high-level tasks such as classification or detection. However, they often fail to capture the intricate details and spatial relationships necessary for precise boundary delineation and detailed feature extraction. This limitation is particularly evident in masked image modeling (MIM) approaches such as MAE [16], SimMIM [47], and iBOT [56], which, while effective at learning global features, struggle to preserve fine-grained details. Consequently, these models often require extensive training epochs to achieve moderate performance improvements, yet still fall short in tasks that demand a high degree of granularity. The lack of textual context in these SSL models further exacerbates their limitations, as they miss the additional semantic guidance that text provides, which is crucial for enhancing the model’s ability to focus on detail-oriented tasks. Incorporating text in the pre-training process can significantly improve the model’s performance by aligning visual features with rich semantic information, leading to better generalization and more accurate granular tasks [2, 49, 52].
We present DetailCLIP, a novel framework designed to overcome the limitations of traditional CLIP-based models and self-supervised learning approaches in tasks involving detail-oriented analysis. DetailCLIP excels in capturing both high-level semantic understanding and fine-grained details, addressing the critical shortcomings of existing models. With its tailored features, DetailCLIP delivers superior performance in scenarios requiring meticulous attention to detail.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
a person crafting pottery on a potter’s wheel, installed in a lowered section of the floor. Two dogs sitting on the grass; one is a golden retriever, and the other is a basset hound.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A black bird with a red beak perched on a tree branch against a blue sky. A large stupa with a golden spire in a mountainous landscape under a blue sky.
Figure 1: This image illustrates the mechanism of token removal. The process selectively eliminates less relevant tokens, retaining only those with strong semantic significance. This approach enhances the model’s focus on the most important aspects of the input, ensuring that critical information is preserved while unnecessary details are discarded.
DetailCLIP introduces an attention-focus mechanism, token removal, that selectively retains the most relevant parts of an image, closely linked to the corresponding detailed-oriented tasks and textual descriptions. As illustrated in Figure 1, this mechanism highlights the critical areas of the image that are crucial for the task, ensuring that the model focuses on the most pertinent details. By integrating textual context, DetailCLIP enhances the model’s ability to understand and emphasize the critical aspects of an image that are most relevant to the task at hand. This targeted attention not only helps capture subtle nuances and intricate details but also ensures that these elements are accurately represented in the model’s output.
DetailCLIP addresses the issue of inadequate pixel-level accuracy by integrating a pixel-level reconstruction feature that emphasizes the preservation and reconstruction of detailed image features. This approach enables the model to effectively handle the intricate information required for fine-grained visual tasks, making it particularly effective in scenarios where pixel-level precision is critical. While other models, such as MAE [16] and SimMIM [47], also incorporate pixel-level reconstruction to enhance accuracy, they typically rely on random token removal, which can miss essential details. Our approach differs by employing a token removal process based on an attention-focus mechanism, ensuring that only the most relevant features are preserved. This targeted strategy allows DetailCLIP to excel in tasks where precise pixel-level accuracy is vital, distinguishing it from previous methods.
Moreover, to enhance detail preservation and overall model robustness, DetailCLIP incorporates a self-distillation strategy [5]. In this approach, the model simultaneously acts as both a student and a teacher: the student model learns by reconstructing missing details from masked images, while the teacher model operates on the original, unmasked images. This iterative learning process allows the model to progressively refine and recover intricate features, resulting in more comprehensive and effective feature extraction. This not only enhances segmentation accuracy but also ensures that DetailCLIP performs exceptionally well in coarse-grained tasks like classification, where maintaining high-level semantic integrity alongside detailed feature recognition is crucial. Previous works, such as SLIP [33], MaskCLIP [12], and A-CLIP [48], have employed self-distillation techniques that primarily focus on coarse-grained comparisons between student and teacher models. In contrast, our approach extends this framework by incorporating both coarse-grained and fine-grained comparisons, enabling the model to capture and preserve intricate details more effectively. This comprehensive strategy enhances DetailCLIP’s performance across a wide range of tasks, making it particularly well-suited for tasks that demand high levels of precision.
Finally, similar to CLIP-based models [38, 31], DetailCLIP emphasizes the crucial connection between vision and language, which is essential for tasks that require a deep understanding of visual content and its associated textual concepts.
To summarize, our contributions are fourfold:
  • We introduce a novel token removal mechanism driven by attention, which selectively retains tokens with strong semantic connections to the corresponding text descriptions along with detailed-oriented tasks. This targeted focus allows the model to emphasize the most relevant regions of an image, significantly enhancing its capability to capture fine-grained details.
  • Our model implements a self-distillation strategy that compares patch-level features using masked images. This approach helps the model to effectively learn low-level semantic features while maintaining a deep understanding of the image content.
  • In addition to patch-level comparisons, our model also employs pixel-level reconstruction of masked images. This technique ensures that the model can accurately reconstruct and preserve intricate visual details, leading to superior performance in tasks that require high precision.
  • By addressing the limitations of existing CLIP-based and traditional self-supervised learning approaches, our model achieves a robust balance between high-level semantic understanding and precise detail extraction. This makes it particularly effective for zero-shot tasks and fine-tuning in applications that demand meticulous attention to detail.
2 Related Work
2.1 Self Supervised Learning
In recent years, Self-Supervised Learning (SSL) has emerged as a dominant force in visual pre-training, driven by its ability to extract meaningful visual features from large-scale unlabeled image datasets [3, 39, 1]. Various pre-training strategies have fueled the success of SSL, each focusing on different pretext tasks. Among these, contrastive learning [35, 7, 41, 17], Masked Image Modeling (MIM) [2, 16, 47, 32, 36], Masked Frequency Modeling (MFM) [46, 29, 53], and self-supervised Knowledge Distillation (KD) [22, 56, 8, 9, 5] have garnered significant attention.
Contrastive learning, a popular SSL technique, aligns different views of the same image within a shared embedding space while distinguishing them from views of other images [35, 7]. Although this approach is effective for learning global features, it often faces challenges in handling tasks that require detailed precision, such as image segmentation. The reason is that contrastive learning primarily focuses on aligning overall image representations, which can lead to the neglect of fine-grained, pixel-level details.
On the other hand, Masked Image Modeling (MIM) has shown greater promise for tasks requiring detailed feature extraction. MIM involves masking portions of an image and training models to reconstruct or predict the missing parts, encouraging the model to understand the intricate relationships between visible and hidden regions [2, 47, 16, 49].
2.2 CLIP-Based Approaches
The integration of self-supervised learning with vision-language pre-training has driven the development of several CLIP-based models, each striving to bridge the gap between visual and textual data [6, 55, 26, 20]. These models build on the original CLIP framework, introducing innovations to enhance the quality of representations and improve alignment between images and text.
For instance, SLIP [33] enhances CLIP by incorporating self-supervised learning with image-to-image contrastive learning, resulting in richer and more robust visual representations. MaskCLIP [12] advances this concept by integrating masked image modeling, which refines visual features by focusing on specific image regions, ensuring they are more closely aligned with the accompanying text. A-CLIP [48] further refines this approach by implementing an attentive token removal strategy, selectively preserving tokens that are semantically relevant to the text, thereby enhancing the precision of the visual-text alignment.
While these methods enhance the original CLIP model by improving representation learning and efficiency, they primarily focus on global alignment between images and text. This emphasis on global features makes them less effective for detail-oriented tasks, where capturing and preserving fine-grained visual details is crucial [21, 44, 19, 58]. In many of these models, token removal is often random or solely based on textual information [12, 48]. In contrast, DetailCLIP addresses these limitations by employing an attention-based mechanism that incorporates both textual information and fine-grained tasks, ensuring better performance in detail-sensitive applications.
3 Method
Figure 2: Overview of the DetailCLIP framework for fine-grained segmentation. The framework uses a teacher-student architecture where the teacher model processes two views of the input image $x$ to generate attention values. These values guide the selective token removal process in the student model, focusing on the most semantically relevant parts of the image. The student model then learns to reconstruct the image through a vision decoder while retaining high-level semantic understanding through a combination of classification token loss ($\mathcal{L}_{CLS}$) patch-level contrastive loss ($\mathcal{L}_{Patch}$), and pixel-level reconstruction loss ($\mathcal{L}_{Rec}$). The model also maintains a strong connection between the vision and text encoders using the CLIP loss ($\mathcal{L}_{CLIP}$), enhancing the model’s segmentation capabilities. In the figure, $g$ is the vision encoder, $h$ is the projection head, $d$ is the image decoder, and $e$ is the text encoder.
3.1 Preliminary and Background
CLIP is a pioneering vision-language model developed by Radford et al. [38]. Its central concept is to utilize a large-scale contrastive learning method to align visual and textual representations. This is accomplished by jointly training an image encoder and a text encoder to increase the similarity between matching image-text pairs while reducing it for mismatched pairs. The training objective of CLIP is based on a symmetric cross-entropy loss applied to cosine similarities between the image and text embeddings. The loss function encourages the correct pairs to have higher similarity than incorrect pairs. More precisely, let $I_{i}$ and $T_{j}$ be the embeddings of the $i$-th image and $j$-th text in the batch, respectively. The cosine similarity between an image and a text is given by:
$s_{i,j}=\frac{I_{i}\cdot T_{j}}{\|I_{i}\|\|T_{j}\|},$ (1)
The logits are scaled by a learnable parameter $\tau$:
$\text{logits}_{i,j}=\frac{s_{i,j}}{\tau},$ (2)
Let $N$ be the number of pairs of image-text in a batch. The loss function for the image-to-text matching is defined as:
$\mathcal{L}_{I2T}=-\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{\exp(\text{logits}_{i,i})}{\sum_{j=1}^{N}\exp(\text{logits}_{i,j})}\right),$ (3)
Similarly, the loss function for the text-to-image matching is:
$\mathcal{L}_{T2I}=-\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{\exp(\text{logits}_{i,i})}{\sum_{j=1}^{N}\exp(\text{logits}_{j,i})}\right),$ (4)
Finally, the overall loss for CLIP is the average of the two losses:
$\mathcal{L}_{CLIP}=\dfrac{1}{2}(\mathcal{L}_{I2T}+\mathcal{L}_{T2I}),$ (5)
3.2 DetailCLIP Framework
As shown in Figure 2, our framework can be outlined through the following steps, which will be discussed in greater detail.
3.2.1 Patch-Level Comparison
We employ a teacher-student framework in which the student model is trained to predict both fine-grained and coarse-grained features of images generated by the teacher model. The teacher model provides ”target” features for the student to learn. For each input, we consider two views: the original views are fed into the teacher model, while their masked versions are used as input for the student model. For further details on the masking process, refer to Section 3.2.3. The student model is tasked with predicting both the masked tokens and the global features generated by the teacher.
Similar to methods used in previous works like DINO [5], the teacher model’s parameters are updated using an exponential moving average (EMA) of the student model’s parameters and its own. Specifically, if $\theta$ represents a parameter of the teacher model and $\phi$ represents a parameter of the student model, the update rule for $\theta$ is $\theta\leftarrow(1-\lambda)\phi+\lambda\theta$, that $\lambda$ starting at 0.996 and gradually increasing to 1.0, following a cosine scheduling approach as outlined in previous works [10, 12, 48]. This strategy ensures that the EMA model remains stable and effective in capturing the most relevant features during training [15].
We employ two primary loss functions to train the student model to ensure both global and fine-grained feature learning are effectively captured.
Global Loss (CLS Token): The global loss, applied to the [CLS] token, measures the difference between the probability distribution of the [CLS] token in the student model and the [CLS] token in the teacher model using KL divergence:
$\mathcal{L}_{\text{CLS}}=\text{D}_{KL}\left(p_{\text{s}}^{\text{[CLS]}}\|p_{\text{t}}^{\text{[CLS]}}\right)$ (6)
Where:
  • $p_{\text{s}}^{\text{[CLS]}}$ is the predicted probability distribution of the [CLS] token from the student model.
  • $p_{\text{t}}^{\text{[CLS]}}$ is the target probability distribution of the [CLS] token from the teacher model.
  • $\text{D}_{KL}$ represents the Kullback-Leibler divergence [24].
Fine-grained Loss (Patch): The patch loss also utilizes KL divergence, comparing the distribution of patch tokens in the student model to those in the teacher model:
$\mathcal{L}_{\text{patch}}=\frac{1}{|\mathcal{M}|}\sum_{j\in\mathcal{M}}\text{D}_{KL}\left(p_{\text{s}}^{j}\|p_{\text{t}}^{j}\right)$ (7)
Where:
  • $\mathcal{M}$ is the set of indices corresponding to the masked patches.
  • $p_{\text{s}}^{j}$ is the predicted probability distribution of the $j$-th masked patch from the student model.
  • $p_{\text{t}}^{j}$ is the target probability distribution of the $j$-th masked patch from the teacher model.
  • $\text{D}_{KL}$ represents the Kullback-Leibler divergence.
3.2.2 Pixel-Level Reconstruction
As illustrated in Figure 2, following token removal, we employ an autoencoding method ($g_{\phi}$-$d_{\phi}$) (MAE  [16]) to reconstruct the original signal from a partially observed input.
By processing only the visible patches, the encoder significantly reduces the amount of data that needs to be handled, thereby improving the model’s efficiency. The decoder is then tasked with reconstructing the original image from the latent representation produced by the encoder. It accomplishes this by using both the encoded visible patches and a set of mask tokens that represent the masked areas. The decoder is smaller and lighter than the encoder, as it is only needed during pre-training. Additionally, by using a smaller decoder, the majority of the computational effort is focused on training the encoder, resulting in a more robust and powerful encoder.
This process is especially well-suited for fine-grained tasks because it pushes the encoder to extract detailed information from the limited visible patches, sharpening its ability to recognize and represent intricate features. By concentrating on these visible patches, the encoder becomes skilled at capturing subtle details that are critical in precision-demanding tasks like segmentation or detailed object recognition. The decoder’s responsibility in reconstructing the masked portions of the image ensures that these fine details are preserved and accurately restored, thereby enhancing the model’s capability to manage complex visual data.
During pre-training, the model learns to reconstruct the original image from the masked input. The loss function ($\mathcal{L}_{\text{Rec}}$) employed during this training is the Mean Squared Error (MSE) between the reconstructed and original images, calculated solely on the masked patches.
$\mathcal{L}_{\text{Rec}}=\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\|\hat{X}_{i}-X_{i}\|^{2}$ (8)
where
  • $\mathcal{M}$ is the indices corresponding to the masked patches.
  • $\hat{X}_{i}$ is the reconstructed patch for the $i$-th masked patch.
  • $X_{i}$ is the original patch for the $i$-th masked patch.
[…]
ADE20K COCO
Methods Dataset Epoch Effective View UperNet Linear AP^b AP^m
Self-Supervised
DeiT[42] IN-1K 300 720M 47.4 - 44.1 39.8
MAE[16] IN-1K 800 960M 46.5 34.3 46.2 39.1
DINO[5] IN-1K 800 9600M 46.8 34.5 47.4 40.1
iBOT[56] IN-1K 300 720M 47.3 34.7 48.4 42.1
AttMask[22] IN-1K 300 4320M 47.5 35.2 48.9 42.2
CLIP-Based Model
CLIP[38] 400M - - 46.4 34.2 43.6 39.5
SLIP[33] YFCC-15M 25 750M 46.6 36.1 44.0 40.3
MaskCLIP[12] YFCC-15M 25 750M 47.5 36.3 45.4 40.9
A-CLIP[48] YFCC-15M 25 750M 47.0 34.7 45.8 41.7
DetailCLIP YFCC-15M 25 750M 48.1 37.3 48.9 42.5
DetailCLIP YFCC-15M 50 1500M 48.8 39.3 50.1 43.3
Table 1: Performance comparison of various models on detail-oriented visual tasks, including segmentation and object detection. All models utilize the vision component’s ViT-B (Vision Transformer Base) architecture. Segmentation results are presented for the ADE20K dataset using UperNet and linear decoders. At the same time, object detection performance is evaluated on the MS COCO dataset with Average Precision for the bounding box. DetailCLIP consistently outperforms other models, particularly in tasks requiring high levels of detail and precision. (”Effective View” is the total number of image views used during training, calculated as the number of views per image multiplied by the number of epochs and size of the dataset. This metric is critical to our evaluation as it reflects the model’s exposure to diverse perspectives of the training data, which directly impacts its ability to generalize and perform well on detail-oriented tasks.)
3.2.3 Token Removal
Token removal in image models, particularly in ViT, is a strategy to refine the attention mechanism by selectively eliminating tokens that contribute less to the final decision-making process. The token removal method allows the model to dynamically down-weight or ignore less important patches, thereby reducing computational load and focusing attention on the most relevant areas of the image, ultimately enhancing both performance and efficiency.
Various approaches for token removal have been proposed, such as random token removal [22] and attentive token removal [48]. In our work, we introduce a novel technique that considers the specific functions of our model, including textual information processing, patch comparison, and image reconstruction. As demonstrated in 4.2, our token removal strategy outperforms existing methods by addressing both fine-grained and coarse-grained details simultaneously.
As illustrated in Figure 2, our token removal process operates as follows. We first input an image view into the teacher encoder model ($g_{\theta}$). The attention values generated by the teacher encoder are then used to mask the $50\%$ of patches with the lowest values. Specifically, for each patch $p_{\text{t}}$ the attention values are calc ulated using the following formula:
$AV_{p_{\text{t}}}=\frac{1}{N}\sum_{i=1}^{N}\text{Softmax}\left(\frac{Q_{i}\cdot K_{i}(p_{\text{t}})}{\sqrt{d}}\right)$ (9)
  • $N$ is the total number of attention heads across all layers.
  • $Q_{i}$ is the query of the [CLS] from the $i$-th attention head.
  • $K_{i}(p_{t})$ is the key of the patch at position $p_{\text{t}}$ from the $i$-th attention head.
  • $d$ is the dimensionality of the query and key.
As shown in Figure 1, masking image patches by removing the lowest half of the attention values (AV) calculated by the teacher encoder results in the images displayed in the figure.
3.2.4 Integrated Loss Function
In this section, we introduce the formulation of a comprehensive loss function that combines the various loss components essential for effective multi-task learning. The integrated loss function is defined as:
$\mathcal{L}_{\text{tot}}=\alpha_{1}\times\mathcal{L}_{\text{CLS}}+\alpha_{2}\times\mathcal{L}_{\text{Patch}}+\alpha_{3}\times\mathcal{L}_{\text{Rec}}+\mathcal{L}_{\text{CLIP}},$ (10)
In this equation, the hyperparameters $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ determine the relative importance of each loss term. These weights are set to $1$ in our experiments unless specified otherwise. The comprehensive loss function is designed to enable the model to concurrently learn across the four tasks introduced in previous sections, ensuring a balanced and integrated training process.
Additionally, in the ablation study, we explore the impact of different hyperparameter choices, demonstrating how varying these weights can influence the overall model performance and task-specific outcomes.
4 Experiments
4.1 Setup
Our computational setup is designed to support the extensive experiments conducted in this study. It consists of four nodes, each equipped with four NVIDIA A100 GPUs, each with 80GB.
4.1.1 Training Data and Augmentation Strategy
Our model is trained on a carefully curated 15 million image subset [38] of the YFCC100M dataset [40], which contains only English-language titles and descriptions. For each image, a valid caption—either a title or description—is randomly selected during training, following the methodology used in SLIP [33].
To enhance model robustness, we employ a data augmentation strategy similar to that of SLIP. Images are randomly resized and cropped to a size between $50\%$ and $100\%$ of their original dimensions. This augmentation is applied to the images in the online training branches, allowing the model to learn from varied perspectives of the same image.
For the teacher part, a slightly different approach is taken. We use a larger randomly cropped sub-image compared to the online views, which allows us to accurately compute attention value.
Flickr30K COCO IN-1K
Methods E I2T T2I I2T T2I 0-Shot
CLIP 25 51.4 32.6 27.9 17.6 37.6
[…]
MaskCLIP 25 60.0 38.8 34.1 21.2 42.7
[…]
CLIP 50 53.9 35.8 30.2 19.2 39.4
[…]
Table 2: Performance comparison of various models on text-image retrieval tasks across the Flickr30K and MS COCO datasets, as well as zero-shot evaluation on ImageNet-1K (IN-1K). The table reports results after 25 and 50 epochs ($E$), showing performance metrics for image-to-text (I2T) and text-to-image (T2I) retrieval on each dataset. The results were collected in the A-CLIP paper [48].
4.1.2 Architecture and Training Setting
Figure 2 illustrates the architecture of the DetailCLIP framework, which is composed of six key components: two visual encoders, denoted as $g_{\theta}$, and $g_{\phi}$, a text encoder named $e_{\gamma}$, a vision decoder named $d_{\phi}$, and two heads referred to as $h_{\theta}$ and $h_{\phi}$.
To ensure a fair comparison with existing models, our visual encoders ($g_{\theta}$ and $g_{\phi}$) are based on the widely recognized Vision Transformer (ViT-B/16) [13] architecture. This model features $12$ layers, each with a width of $768$ and $12$ attention heads.
For the text encoder ($e_{\gamma}$), we adopted a 12-layer Transformer architecture with a width of $512$ and $8$ attention heads, following the design principles established by the CLIP model [38].
The decoder ($d_{\phi}$) itself is composed of a series of transformer blocks like MAE paper [16]. The decoder handles the entire set of input patches, including both the attention-masked and retained patches. Positional embeddings are added to each token in the set to preserve the spatial information.
The projection heads ($h_{\theta}$ and $h_{\phi}$) in our framework are implemented as 3-layer MLPs, featuring an L2-normalized bottleneck, similar to the approach used in DINO [5]. The output dimension of the shared projection head is set to $8192$ like iBOT paper [56], for robust feature representation and alignment during training.
For our experiments, we employ the $AdamW$ optimizer [30] with a learning rate of $45e-$ 4 and a weight decay of $0.5$. The training is conducted with a substantial batch size of $4096$.
4.2 Experimental Analysis
Epochs
Methods
SUN397
Aircraft
DTD
Pets
Flowers
GTSRB
MNIST
STL-10
KITTI
Country211
PCAM
CLEVR
SST2
Average
[…]
Table 3: Zero-shot performance evaluation across multiple classification benchmarks. Epochs indicate the total number of training epochs. The results were collected in the A-CLIP paper [48].
4.2.1 Detail-Oriented Visual Tasks
In order to comprehensively evaluate the effectiveness of our proposed DetailCLIP framework, we designed a series of experiments focused on detail-oriented visual tasks. Specifically, we implemented different distinct tasks that emphasize the model’s ability to capture fine-grained details in complex visual environments.
Semantic Segmentation on ADE20K: To evaluate the segmentation capabilities of DetailCLIP, we conducted a series of experiments on the ADE20K dataset [54] using different decoding strategies. First, we tested the model with the UperNet [45] decoder, a well-established architecture for semantic segmentation, to assess its ability to delineate object boundaries and achieve high segmentation accuracy. Next, we employed a linear decoder for segmentation tasks to explore the model’s versatility. This more straightforward approach allowed us to evaluate how well DetailCLIP maintains performance with minimal architectural complexity. These experiments collectively provide a comprehensive understanding of DetailCLIP’s strengths in detail-oriented segmentation tasks.
We utilized the UperNet or linear decoder with an input resolution of $512\times 512$ pixels, conducting end-to-end training for $160k$ iterations. The performance of our model was evaluated using the mean Intersection over Union (mIoU) metric.
Object Detection and Instance Segmentation on the COCO: This task requires simultaneous object localization and classification, making them complex tasks that demand high precision. For our evaluation, we employed the Cascade Mask R-CNN model[4, 18] like iBOT [56], known for effectively producing bounding boxes and instance masks on the COCO dataset [27]. The evaluation metrics used in this study were the Average Precision for bounding boxes ($AP^{b}$) and masks ($AP^{m}$).
Evaluating Task Outcomes: Table 1 highlights the significant advantages of DetailCLIP over existing models in various detail-oriented visual tasks. Utilizing the ViT-B architecture, DetailCLIP outperformed other models across multiple benchmarks. For instance, on the ADE20K dataset with the UperNet decoder, DetailCLIP achieved a $mIoU$ of $48.8$, $1.3$ points higher than the closest competitor, MaskCLIP, which recorded a $mIoU$ of $47.5$. Similarly, in the linear decoder setting, DetailCLIP’s $mIoU$ of $39.3$ surpasses SLIP’s performance by $3.2$ points, highlighting our model’s superior capacity to maintain high performance.
Furthermore, in object detection tasks on the MS COCO dataset, DetailCLIP for bounding boxes achieved $48.9$ $AP^{b}$, outperforming previous models by as much as $3.1$ points, with ACLIP being the closest at $45.8$. The mask precision is $42.5$ $AP^{m}$, further underscoring our model’s effectiveness, exceeding the best-performing alternative by $0.8$ points. These substantial improvements demonstrate DetailCLIP’s effectiveness in enhancing high-level semantic understanding and fine-grained detail extraction, setting a new benchmark in the field.
Visual Comparative Analysis: Figure 5 shows the result of our model compared to the other baselines. We invite you to refer to the appendix. This comparison covers segmentation and object detection tasks, demonstrating DetailCLIP’s enhanced performance.
Image Ground Truth A-Clip SLIP DetailCLIP
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Visual comparison of segmentation results using the Linear decoder across different models. The comparison includes SLIP, Attentive Mask CLIP, and DetailCLIP at 25 epochs.
4.2.2 Image Classification
Zero-shot on Text-Image Retrieval: We also present the zero-shot text-image retrieval on 3 benchmark datasets: Flickr30K [50], MS-COCO [27], and ImageNet-1K [11]. Our findings indicate that using plain text, without any added prefixes or suffixes, consistently yields strong performance across all models evaluated.
Table 2 comprehensively evaluates different models on text-image retrieval tasks, explicitly focusing on Flickr30K and MS-COCO datasets, also a zero-shot evaluation on ImageNet-1K.
At 25 epochs, DetailCLIP demonstrates the highest performance across most metrics, achieving the best results in Flickr I2T (62.8), T2I (42.2), COCO I2T (38.3), and zero-shot on IN-1K (43.9), closely matching A-CLIP. For 50 epochs, DetailCLIP maintains its superior performance, particularly in the COCO dataset, where it ties in with A-CLIP for the best I2T score (39.8) and surpasses A-CLIP in T2I with a score of 24.9. The results suggest that DetailCLIP is particularly strong in tasks that require detailed matching between text and images. DetailCLIP outperforms other models in both datasets’ I2T and T2I retrieval tasks, indicating its effectiveness in text-image association.
Zero-Shot Classification Performance Evaluation on Diverse Benchmarks: We evaluated the zero-shot classification performance of our proposed DetailCLIP approach across $13$ different classification tasks, following the evaluation protocols established by SLIP [33]. This included using identical prompts during testing to ensure a fair comparison. This rigorous evaluation framework allowed us to directly assess and compare DetailCLIP’s performance against other leading models.
The results in the table 3 demonstrate that DetailCLIP consistently outperforms its competitors. Specifically, DetailCLIP achieved the highest performance in $12$ out of $26$ evaluation scenarios and secured the second-best position in $5$ additional cases. Notably, DetailCLIP also achieved the highest average scores in both the 25-epoch ($37.3$) and 50-epoch ($37.6$) settings, underscoring its robustness and generalization capabilities across a diverse array of benchmarks.
These results highlight DetailCLIP’s effectiveness in zero-shot classification tasks, particularly compared to other state-of-the-art models across various datasets.
[…]
Table 4: Impact of varying $\alpha$ values on the zero-shot accuracy of ImageNet-1k after 25 epochs of training with the DetailCLIP.
4.2.3 Ablation Study on Loss Weight Optimization
In this ablation study, we investigate the impact of different weight configurations on the comprehensive loss function $\mathcal{L}_{\text{tot}}$ introduced in Eq. 10. We aim to identify the optimal values for the hyperparameters $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ that effectively balance the various losses.
Table 4 presents the results of this exploration, showing how different combinations of these weights influence the zero-shot accuracy on the ImageNet-1k dataset after $25$ epochs of training. The accuracy values reflect the model’s performance under varying degrees of emphasis on each loss term.
Analyzing the results, we find that the baseline configuration, where all weights are set to 1, achieves the highest accuracy of $43.9\%$, suggesting that equal emphasis on all loss components provides a robust performance baseline. When $\alpha_{3}$ is reduced to 0, thereby ignoring the reconstruction loss, the accuracy slightly decreases to $42.9\%$, indicating that while this loss component is not critical, it still positively influences overall model performance. Doubling $\alpha_{3}$ to 2 results in a marginal accuracy decrease to $43.2\%$, implying that an overemphasis on the reconstruction loss can detract slightly from the model’s effectiveness in other tasks. Lowering the weights of $\alpha_{1}$ and $\alpha_{2}$ to $0.5$ yields an accuracy of $43.3\%$, demonstrating that moderate weight reductions do not significantly impact performance. However, eliminating these ($\alpha_{1}=0$ and $\alpha_{2}=0$) reduces accuracy further to $42.6\%$, underscoring their essential role in the model’s learning process.
5 Conclusion
In this paper, we introduced DetailCLIP, a novel framework designed to overcome traditional CLIP-based models’ limitations in fine-grained and coarse-grained tasks. While existing models like CLIP excel at globally aligning image and text representations, they often fail to capture the fine-grained details necessary for tasks such as image segmentation and object detection. To address this, we integrated innovative techniques into DetailCLIP, including patch-level comparison, pixel-level reconstruction, and an attention-based token removal mechanism. These features enable the model to focus on the most critical regions of an image, significantly enhancing its performance in detail-oriented tasks. Our extensive experiments demonstrated that DetailCLIP consistently outperforms state-of-the-art models in tasks requiring high levels of detail, such as segmentation on the ADE20K dataset and object detection on the MS COCO dataset, where it sets new benchmarks in critical metrics like mIoU and Average Precision. Furthermore, DetailCLIP’s strength extends beyond fine-grained tasks, as it excels in coarse-grained tasks like zero-shot classification.
Appendix A Visual Analysis
This section presents a comprehensive comparative analysis of DetailCLIP’s performance against several baseline models, primarily focusing on segmentation and object detection tasks. Through a series of detailed visualizations, we provide side-by-side comparisons that emphasize the superior capability of DetailCLIP in capturing and reconstructing fine-grained details, which is crucial for achieving high segmentation accuracy. Specifically, Figure 4 showcases the semantic segmentation results using the UPerNet decoder, while Figure 5 highlights the results with a linear decoder. These figures emphasize the notable improvements achieved by DetailCLIP in handling complex images, particularly when it comes to accurately delineating boundaries and preserving detailed features, areas where baseline models often struggle. The segmentation comparisons are conducted using the ADE20K dataset.
In addition to segmentation, Figures 6 and 7 present the object detection results using the MS COCO dataset, further demonstrating the effectiveness of DetailCLIP across multiple vision tasks. By visualizing these outputs, we showcase DetailCLIP’s enhanced performance and shed light on the limitations of other models, such as SLIP and A-CLIP. These comparisons underscore the significant advancements made by DetailCLIP in tackling detail-oriented tasks, solidifying its robustness and versatility in a wide array of visual challenges.
Overall, these visual and quantitative evaluations illustrate the significant advancements made by DetailCLIP in addressing detail-oriented challenges across both segmentation and object detection. The side-by-side comparisons solidify DetailCLIP’s robustness and adaptability, making it a more effective and versatile solution for a wide range of fine-grained vision tasks.
Image Ground Truth A-Clip SLIP DetailCLIP-25 DetailCLIP-50
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Visual comparison of segmentation results using the UPerNet decoder across different models. The comparison includes SLIP, A-CLIP, DetailCLIP at 25 epochs, and DetailCLIP at 50 epochs.
Image Ground Truth A-Clip SLIP DetailCLIP
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Visual comparison of segmentation results using the Linear decoder across different models. The comparison includes SLIP, A-CLIP, and DetailCLIP at 25 epochs.
Image + Objects Image + Objects
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: Object detection results using DetailCLIP after 50 training epochs. The image highlights DetailCLIP’s ability to detect objects accurately, showcasing its precision and effectiveness in complex scenes
Image + Objects Image + Objects
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: Object detection results using DetailCLIP after 50 training epochs. The image highlights DetailCLIP’s ability to detect objects accurately, showcasing its precision and effectiveness in complex scenes